OpenStudy (calculusxy):

Write in point-slope form a line that is parallel to \(-4x + 6y = 36\) passing thru \(2, 3\)

Nnesha (nnesha):

what is the slope of this line -4x+6y=36 ?

OpenStudy (calculusxy):

I got \(\large \frac{2}{3}\) but I am pretty sure that's wrong.

OpenStudy (anonymous):

must be a math teacher made up this question \[-4x+6y=36\] is the same as \[-2x+3y=18\]

OpenStudy (anonymous):

and yes, it is \(\frac{2}{3}\)

OpenStudy (calculusxy):

I know that parallel lines have the same slope?

Nnesha (nnesha):

looks good. y=mx+b is a slope intercept form where m is slope and b is y-intercept and parallel lines have the same slope

Nnesha (nnesha):

correct. \[\huge\rm y-y_1=m(x-x_1)\] point slope from m=slope and passes through (x_1,y_1) point

OpenStudy (calculusxy):

I don't know why, but my paper says that the formula for the point-slope from is \[\large y_1 - y_2 = m(x_1 - x_2) \]

Nnesha (nnesha):

that's the same thing. then it's passes through (x_2 ,y_2)

OpenStudy (calculusxy):

Okay so do I make up another coordinate by using the slope?

Nnesha (nnesha):

no to write it in point slope form you just one coordinate and slope that's it

Nnesha (nnesha):

need**

OpenStudy (calculusxy):

How's that possible can you show me?

Nnesha (nnesha):

it should look like this \[\rm y-9=3(x-34)\]

OpenStudy (calculusxy):

Where did you get those numbers from?

Nnesha (nnesha):

we should keep the x_1 and y_1 and replace 2nd (x_2 ,y_2) with the given point

Nnesha (nnesha):

that's an example .. :D

OpenStudy (calculusxy):

oh okay

OpenStudy (calculusxy):

So would we have : \[\large y - 3 = \frac{ 2 }{ 3 }(x - 2)\]

Nnesha (nnesha):

perfect!

OpenStudy (calculusxy):

Thank you!

Nnesha (nnesha):

yw

OpenStudy (calculusxy):

Also I have another question. If we have a perpendicular slope then it means that we take the slope of the current line and find the negative reciprocal?

OpenStudy (calculusxy):

\[-3 \rightarrow \frac{ 1 }{ 3 }\] Would this work if we have a negative slope?

Nnesha (nnesha):

remember if one of the coordinate is negative lets say we have (-3,-4) then \[y-(-4)=4(x-(-3)) \large\rm \rightarrow y+4=4(x+3)\] you should multiply by the negative sign in the formula

Nnesha (nnesha):

that's correct!

OpenStudy (calculusxy):

Thank you SOOOO MUCH! You're the best as always :)

Nnesha (nnesha):

glad to help! and thanks! o^_^o